OpenStudy (anonymous):

Solvetheproportion: 16/17=x/40 42.5 37.6 -42.5 -37.6

OpenStudy (aroub):



OpenStudy (anonymous):

yes like that

OpenStudy (anonymous):

(16/17)x40 = x x =37.647 (to 3.d.p)

OpenStudy (anonymous):

37.6

OpenStudy (aroub):

you croos multiply.. 16(40)=17(x) 640=17x x=640/17